OpenStudy (anonymous):

Which of the following correctly models and gives the quotient of 1 over 2 ÷ 3 over 10?

OpenStudy (anonymous):

\[\frac{1}{2}\times \frac{10}{3}\]

OpenStudy (anonymous):

cancel first, multiply last, get \[\frac{5}{3}=1\tfrac{2}{3}\]

OpenStudy (anonymous):

ok 1 more question

OpenStudy (anonymous):

Joan represents the height at which a sea gull is flying at +4 feet and the depth at which a fish is swimming at −1 feet on a number line, where 0 represents the surface of the water. Which of these explains whether the sea gull or the fish is closer to the surface of the water?

OpenStudy (anonymous):

as for the model, this looks like something made up the the criminal master minds at FLVS so i have no idea what the model is supposed to be

OpenStudy (anonymous):

i guess the fish is closer because it is one foot away right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

go with C